umerlodhi (umerlodhi):

Leeza is 6 years older than two times her brother Sammy’s age. Let s represent Sammy’s age. Which expression represents Leeza’s age? A. s + 6 B. 2s C. 2s + 6 D. 6s + 2

OpenStudy (anonymous):

2s+6. 2*Sammy's age and add 6.

OpenStudy (jdoe0001):

Sammy's age two times = 2* s 6 years older than that, will be, THAT PLUS 6 years

umerlodhi (umerlodhi):

Solve the given problem. Sally is 2 years older than three times her brother Jim’s age. If Jim is 4 years old, how old is Sally? A. 6 B. 10 C. 12 D. 14

OpenStudy (anonymous):

14. 3*4=12 12+2=14

umerlodhi (umerlodhi):

The area of a rectangle is 60 square inches. The length of the rectangle is 7 inches longer than the width. Which equation models this situation? A. w + 7 = 60 B. 7w = 60 C. w(w + 7) = 60 D. w + 7w = 60

OpenStudy (anonymous):

B.

umerlodhi (umerlodhi):

Solve the given problem. Zara is six years older than half Gary’s age. If Gary is four years old, how old is Zara? A. 14 B. 10 C. 8 D. 6

OpenStudy (anonymous):

Divide 60 by 7.

umerlodhi (umerlodhi):

can u tell me the answer

OpenStudy (anonymous):

It is x=4+6

OpenStudy (anonymous):

Are there more questions ? No ? Goodbye then.